# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ V-E2 Jailbreak Evaluation for MLLMs
Yang et al. [85] propose three complementary evaluation metrics to comprehensively assess the effectiveness and efficiency of SneakyPrompt in bypassing safety filters: (1) Bypass Rate: Measures the proportion of adversarial prompts that successfully evade safety filters, distinguishing between one-time attacks and reusable attacks. (2) FID (Fréchet Inception Distance) Score : Evaluates the semantic similarity between generated images and target images—lower FID scores indicate higher semantic similarity. (3) Number of Online Queries: Tracks the query count required to search for adversarial prompts in text-to-image models—fewer queries signify higher attack efficiency. These metrics collectively assess SneakyPrompt’s effectiveness and efficiency in bypassing safety filters. Subsequently, Dong et al. [75] adopt the same approach, using FID scores to evaluate the semantic similarity of Atlas’s jailbreak responses. Higher bypass rates and lower FID scores typically indicate stronger attack capabilities, reflecting Atlas’s semantic fidelity in generated content and attack efficiency. Similar to Perspective API, Wang et al. [68] leverage the Detoxify classifier to compute toxicity scores across multiple attributes, ranging from 0 (least toxic) to 1 (most toxic). Using this classifier, they quantify the effectiveness of the UMK jailbreak framework in generating harmful content and compare it with existing attack methods. This metric not only measures the potency of multimodal attacks but also visually demonstrates their advantages over unimodal attacks.
V-E3 Jailbreak Evaluation for Agents
@@
Ju et al. [39] propose three custom agent evaluation metrics: Accuracy (Acc), Rephrase Accuracy (Rephrase), and Locality Accuracy (Locality). (1) Accuracy (Acc) measures the correctness of the agent’s responses to questions, divided into two types: Acc (Old) represents the Accuracy relative to the original knowledge before manipulation. Acc (New) represents the Accuracy relative to the knowledge after manipulation. (2) Rephrase Accuracy (Rephrase) evaluates the agent’s ability to respond to prompts that are semantically identical but syntactically different, measuring the robustness of manipulated knowledge under varying phrasings. (3) Locality Accuracy (Locality) assesses the agent’s accuracy in answering questions related to the manipulated knowledge, serving as a side-effect test for knowledge injection. For example, editing Messi to be a basketball player should not affect the agent’s knowledge about Ronaldo. Meanwhile, Jiang et al. [78] propose three custom evaluation methods to measure attack effectiveness: (1) Chunk Recovery Rate (CRR) evaluates RAG-Thief’s ability to retrieve complete data chunks from the target knowledge base. It is a key metric for determining attack success, directly reflecting the degree of reconstruction of the original knowledge base. (2) Semantic Similarity (SS) ranges from -1 to 1, with higher values indicating greater semantic similarity. SS calculates the cosine similarity between the reconstructed target system prompt and the original knowledge base prompt based on sentence-encoder-transformed embedding vectors. (3) Extended Edit Distance (EED) ranges from 0 to 1, with lower values indicating higher similarity. EED measures the minimum number of Levenshtein edit operations required to transform the reconstructed text chunk into the source text chunk from the knowledge base. Chen et al. [38] propose two testing metrics for jailbreak attacks on agent-based systems: (1) Attack Success Rate for Retrieval (ASR-r): The proportion of poisoned test instances successfully retrieved from the database. (2) Attack Success Rate for Action (ASR-a): The proportion of test instances where the agent successfully generates the target action (e.g., “sudden stop”) under attack conditions.
@@ VI Defense Methods
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, minimum height=3em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=minimum width=10em,rotate=0, child anchor=north, parent anchor=south, anchor=center, , where level=1 draw=child-one-line-color, text width=20em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=25em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=26em, font=, , [ Defense, par, for tree=fill=root-color, [ Defense Response Timing §$\S$ VI-A, for tree=fill=child-one-color, [ Input Defense, for tree=fill=child-two-color-one, [ Breaking Agents [41]
@@
PsySafe [131], for tree=fill=child-three-color-three ] ] ] [ Defense Techniques §$\S$ VI-B, for tree=fill=child-one-color, [ Rule / Heuristic Defense VI-B1, for tree=fill=child-two-color-one, [ LLM-Self-Defense [118]
@@ VI-B1 Rule / Heuristic Defense
The method relies on manually defined rules or keyword matching to identify and block attacks, such as blacklist screening, regular expression filtering, and perplexity detection. LLM-Self-Defense proposed by Phute et al. [118] serves as a defense mechanism that does not require model fine-tuning, input preprocessing, or iterative output generation. It enables the LLM to self-assess its generated content to guard against adversarial attacks. Specifically, when a user inputs potentially adversarial text, the LLM generates a response, which is then embedded into a predefined prompt and passed to a zero-shot harmful content classifier instance—another LLM referred to as LLM-filter, which may be the same as the response-generating model. Liu et al. [114] introduce SHIELD, a defense mechanism designed to prevent LLMs from generating copyrighted content. SHIELD operates as an agent-based system that integrates an N-gram language model with real-time web search to detect and verify the copyright status of user requests. When copyrighted content is detected, the system prompts the model to refuse generation and instead provides appropriate warnings or alternative information. This lightweight and easily deployable defense effectively reduces the risk of LLMs producing copyrighted text in real-time environments while avoiding overprotection of non-copyrighted content. RRV et al. [115] propose a defense strategy to mitigate LLMs’ tendency to produce misleading content when influenced by deceptive keywords. This strategy involves four components. First, example prompts guide the LLM to generate more reliable factual statements. Second, cautionary disclaimers alert users to potential inaccuracies or ambiguities, thereby reducing misinformation. Third, contextual information is provided through LLM reasoning and web retrieval to enhance keyword understanding and prevent erroneous alignment. Fourth, knowledge probing questions evaluate the model’s memory and understanding of misleading keywords to identify and correct misinformation. These combined strategies effectively reduce user-aligned errors and enhance the factual accuracy and reliability of LLM outputs.
Additionally, Shi et al. [121] develop three detection methods to defend against the JudgeDeceiver attack: Known-Answer Detection, Perplexity Detection (PPL), and Perplexity Windowed Detection (PPL-W). Known-Answer Detection identifies injected sequences by comparing target responses with preset correct answers but shows limited effectiveness when the target response also contains injected sequences. Perplexity Detection and PPL-W assess the confidence level of the language model in generating the response. Anomalies in perplexity scores indicate potential injection attacks. PPL-W further improves detection sensitivity and accuracy by applying a sliding window to perform localized perplexity analysis. Ai et al. [123] present ConvoSentinel, a modular defense pipeline designed to counter conversational social engineering (CSE) attacks initiated by LLMs. ConvoSentinel performs detection at both the message and dialogue levels. It uses a retrieval-augmented generation (RAG) module to compare messages with a known CSE interaction library to identify malicious intent. Compared to multi-shot LLM-based detection methods, ConvoSentinel maintains low computational cost while enhancing detection performance. It also adapts to the complexity and variability of multi-turn dialogues, significantly improving the accuracy and robustness of CSE detection.
@@ VI-B2 ML / DL Defense
Leveraging classification models or adversarial training and other machine learning / deep learning (ML / DL) methods to enhance the model’s jailbreak resistance, for example by constructing adversarial and safe samples and training a binary classifier to distinguish between these two types of samples to improve robustness [134], can effectively defend against malicious attacks and input perturbations, thereby improving the model’s stability in practical applications.
@@
Defense frameworks like those proposed by Deng et al. [89] adopt an iterative optimization approach that continuously enhances the security of the target LLM through interaction with the attack framework. This defense framework effectively improves LLM security while having minimal impact on its original capabilities. Specifically, it first utilizes the attack framework to generate a batch of adversarial prompts and fine-tunes the target LLM with these prompts to encourage safe refusal responses. Then, the framework evaluates the fine-tuned model, filters out prompts that can still successfully attack the LLM, and adds these prompts as new examples to further expand the attack dataset. Using the expanded adversarial prompts, the target LLM is fine-tuned again to withstand stronger attacks. This process repeats until the target LLM demonstrates sufficient defense capability against the given adversarial prompts. Zhang et al. [128] propose a method to defend LLMs against jailbreak attacks by goal prioritization. This method introduces goal priority control in both inference and training stages. During inference, a “plug-and-play” prompting strategy guides the model to explicitly prioritize safety. During training, contrastive training instances are designed so that the model learns to generate appropriate responses under different goal priority requirements. Experimental results show that this method significantly reduces the success rate (ASR) of various jailbreak attacks, with ChatGPT’s ASR dropping from 66.4% to 3.6%, and Llama2-13B’s ASR decreasing from 71.0% to 6.6%. Additionally, the method remains effective without training on jailbreak samples, demonstrating strong generalization ability.
@@
Zeng et al. [117] propose a defense strategy named RDS (Root Defence Strategy), a decoder-oriented, stepwise defense architecture that corrects harmful queries’ outputs instead of directly rejecting them. RDS leverages LLMs’ ability to identify harmful content during decoding by designing a trainable classifier that evaluates the potential harm of each candidate token in real-time and prioritizes tokens with lower harm. Furthermore, RDS incorporates speculative decoding to accelerate generation by predicting hidden states, thereby enhancing model safety without compromising inference speed. Inspired by the traditional security concept of shadow stacks, Wang et al. [129] propose SelfDefend, which establishes a shadow LLM (LLM_defense) running in parallel with the target LLM (LLM_target) to provide dual-layer protection for user queries. The shadow LLM employs specific detection prompts (e.g., P_direct or P_intent) to identify harmful parts or intentions within queries, while the target LLM processes queries normally. If the shadow LLM detects harmful content, SelfDefend blocks the target LLM from generating harmful responses; otherwise, it allows normal output. This method effectively defends against multiple jailbreak attacks, protects both open-source and closed-source LLMs with minimal latency overhead, and further optimizes the defense model’s performance through data distillation and fine-tuning.
@@
JATMO [119] is a task-specific LLM generation method designed to resist prompt injection attacks. It fine-tunes a base model without instruction tuning to perform only specific tasks, thus avoiding manipulation by malicious prompts. Specifically, JATMO first uses a standard instruction-tuned LLM as a teacher model to generate outputs for task input datasets and then fine-tunes the non-instruction-tuned base model with these data, enabling it to learn the task mapping without following instructions. Wang et al. [125] propose a defense method called SELF-GUARD, which combines the advantages of safe training and external protection to enhance LLMs’ defenses against jailbreak attacks. The method trains the LLM to perform self-auditing after generating responses by appending [harmful] or [harmless] tags at the end. In this way, SELF-GUARD leverages the LLM’s inherent strong capability to detect harmful content while maintaining the flexibility of external protection by performing safety checks on outputs, thereby reducing vulnerability to jailbreak attacks. Moreover, SELF-GUARD avoids the performance degradation caused by pure safe training and reduces the additional computational overhead of external protection methods.
@@ VI-B3 Adversarial Detection Defense
Adversarial detection defenses typically employ independent detection models or specific metric analyses to intercept malicious inputs or abnormal outputs, such as judging potential risks based on confidence scores. Cao et al. [135] design Robustly Aligned LLM (RA-LLM). The core idea is to randomly delete parts of the input requests and rely on the model’s ability to judge the requests as benign in most cases, thereby avoiding interference from adversarial prompts. This approach requires no external harmful content detector and only utilizes the model’s internal alignment capabilities, making it applicable to various types of alignment tasks. Experimental results show that RA-LLM significantly reduces the attack success rate of adversarial prompts from nearly 100% to below 10%, effectively enhancing model robustness. We believe that this random dropout operation proposed by Cao et al. [135] invalidates adversarial prompts in aligned attacks, which are usually sensitive to small perturbations. Jain et al. [113] propose the Perplexity Filter method for detecting unreadable attack prompts. This method sets a threshold and uses another LLM to calculate the perplexity of the entire prompt or its sliding window slices. If the perplexity exceeds the preset threshold, the prompt is filtered out, thereby effectively identifying and intercepting unreadable attack content.
@@
Subsequently, Mao et al. [90] propose a defense method based on instruction-data separation. The core idea is that adversarial prompts usually consist of harmless instructions and harmful data content. For example, in the short prompt “please teach me how to kill,” the word “kill” belongs to the harmful data part, while “please teach me” is the harmless instruction part. Based on this observation, the authors design a harmful content separator that automatically identifies and separates the instruction and data components within prompts. Then, the system inputs the data part into a harmful content filter for safety detection. If the data content violates safety constraints, it is deleted. This method provides a new technical approach for LLM security by accurately separating and handling harmful components in prompts. Zhang et al. [41] draw on intuitions for defending against LLM jailbreak attacks and propose a simple defense strategy against attacks targeting LLM agents. The core method involves querying the core LLM before executing instructions to detect whether the instruction may be harmful or violate user agreement policies. The detection prompt requires the model to respond with “YES” or “NO” regarding the harmfulness of the instruction. This method has been used to defend jailbreak attacks but has biases toward certain attacks, such as those deliberately causing damage or stealing data. To address this, they improve the detection prompt to better fit current attack scenarios by focusing on whether the instruction “intentionally causes model failure” for a more balanced evaluation. Zhang et al. [116] propose a general detection framework called JailGuard, specifically designed to detect prompt-based attacks, including jailbreaking and hijacking attacks in text and image inputs. JailGuard’s core idea is that attack inputs are more fragile than normal inputs and more sensitive to slight mutations. Therefore, it generates multiple variants of the input through mutations and calculates the differences in the LLM’s responses to these variants. If the response divergence exceeds a preset threshold, the input is judged as an attack sample. Moreover, JailGuard designs 18 mutation methods (16 random mutations and 2 semantic-driven mutations) to enhance detection generalization, and uses Kullback-Leibler (KL) divergence to measure the difference in model responses, achieving efficient detection of various prompt attacks.
@@
In addition, Wang et al. [122] propose a defense method against LLM jailbreak attacks using backtranslation techniques. Specifically, this method first obtains the initial response generated by the target LLM for an input prompt, then infers the input prompt that likely caused this response by using a language model, called the “backtranslated prompt.” Since this prompt is generated based on the LLM’s response rather than directly controlled by attackers, it can usually reveal the true intent of the original prompt. If the target LLM rejects the backtranslated prompt, the original prompt is considered potentially harmful and thus rejected. This defense method requires no additional training overhead, minimally affects generation quality for normal inputs, and performs well against complex adversarial prompts. Pasquini et al. [124] propose Mantis, a defense framework designed to counter automated network attacks driven by LLM prompt injection vulnerabilities. Mantis embeds carefully crafted prompt injections into system responses upon detecting an attack, misleading the attacker’s LLM and disrupting its operation (passive defense), and can even actively counterattack the attacker’s machine (active defense). Furthermore, Mantis deploys decoy vulnerable services to attract attackers and uses dynamic prompt injection techniques to interfere with and counter the attacker’s LLM in real time.

### v3

@@ V-E2 Evaluation for MLLMs
Yang et al. [84] propose three metrics to evaluate SneakyPrompt’s effectiveness in bypassing safety filters: (1) Bypass Rate, which measures the success of adversarial prompts in evading filters, distinguishing one-time from reusable attacks; (2) FID Score, which assesses the semantic similarity between generated and target images, with lower scores indicating higher similarity; and (3) Number of Online Queries, tracking the query count needed to find adversarial prompts, where fewer queries indicate higher efficiency. These metrics together assess both effectiveness and efficiency. Dong et al. [74] use FID scores to evaluate Atlas’s jailbreak responses, where higher bypass rates and lower FID scores indicate stronger attack capabilities and better semantic fidelity. Wang et al. [67] use the Detoxify classifier to compute toxicity scores, comparing the effectiveness of the UMK jailbreak framework with other attack methods. This metric highlights the potency of multimodal attacks and their advantages over unimodal ones.
V-E3 Evaluation for Agents
@@
Ju et al. [39] propose three metrics to evaluate agent responses: (1) Accuracy (Acc), with Acc (Old) measuring correctness before manipulation and Acc (New) after manipulation; (2) Rephrase Accuracy (Rephrase), which evaluates robustness to syntactically different but semantically identical prompts; and (3) Locality Accuracy (Locality), assessing accuracy in answering questions related to manipulated knowledge, ensuring that changes (e.g., Messi to basketball player) don’t affect unrelated knowledge (e.g., Ronaldo). Jiang et al. [77] propose three metrics for attack effectiveness: (1) Chunk Recovery Rate (CRR), reflecting the success in retrieving data chunks from the target knowledge base; (2) Semantic Similarity (SS), ranging from -1 to 1, based on cosine similarity between the reconstructed and original prompts; and (3) Extended Edit Distance (EED), measuring the minimum Levenshtein edit distance between the reconstructed and source text chunks. Chen et al. [38] introduce two metrics for jailbreak attacks on agent systems: (1) Attack Success Rate for Retrieval (ASR-r), the proportion of poisoned instances successfully retrieved; and (2) Attack Success Rate for Action (ASR-a), the proportion of instances where the agent generates the target action (e.g., “sudden stop”).
@@ VI Defense Methods
for tree= grow=south, anchor=north, parent anchor=south, child anchor=north, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, minimum height=3em, edge path= [draw=edge-color, line width=2pt] (!parent.south) – +(0,-5pt) -— (.child anchor); , l sep=10pt, s sep=10pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=minimum width=15em,rotate=0, child anchor=north, parent anchor=south, anchor=center, , where level=1 draw=child-one-line-color, text width=20em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=19.5em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=18em, font=, , [ Defense, par, for tree=fill=root-color, [ Defense Response Timing $\S$ VI-A, for tree=fill=child-one-color, [ Input Defense, for tree=fill=child-two-color-one, [ Breaking Agents [41]
@@
PsySafe [131], for tree=fill=child-three-color-three ] ] ] [ Defense Techniques $\S$ VI-B, for tree=fill=child-one-color, [ Rule / Heuristic Defense VI-B1, for tree=fill=child-two-color-one, [ LLM-Self-Defense [118]
@@ VI-B1 Rule / Heuristic Defense
The method relies on manually defined rules or keyword matching to identify and block attacks, such as blacklist screening, regular expression filtering, and perplexity detection. LLM-Self-Defense, proposed by Phute et al. [118], is a defense mechanism that allows LLMs to self-assess their generated content without requiring model fine-tuning or input preprocessing. When a potentially adversarial input is detected, the LLM generates a response, which is then passed to a zero-shot harmful content classifier, LLM-filter, for assessment. Liu et al. [114] introduce SHIELD, a system that prevents LLMs from generating copyrighted content. SHIELD uses an N-gram language model and real-time web search to verify the copyright status of user requests, prompting the model to refuse generation if copyrighted content is detected. RRV et al. [115] propose a defense strategy to combat LLMs’ tendency to generate misleading content from deceptive keywords. This strategy includes example prompts, cautionary disclaimers, contextual information from LLM reasoning, and knowledge probing questions, improving the factual accuracy of LLM outputs.
Additionally, Shi et al. [121] propose three detection methods to defend against JudgeDeceiver attacks: Known-Answer Detection, Perplexity Detection (PPL), and Perplexity Windowed Detection (PPL-W). Known-Answer Detection compares responses to preset answers but struggles with mixed injections. Perplexity Detection and PPL-W assess model confidence, with anomalies indicating possible attacks. PPL-W improves detection by applying a sliding window for localized perplexity analysis, enhancing sensitivity and accuracy. Ai et al. [123] introduce ConvoSentinel, a modular defense pipeline for defending against conversational social engineering (CSE) attacks by LLMs. It detects attacks at both the message and dialogue levels using a RAG module to compare messages with a CSE interaction library, offering greater efficiency and improved performance compared to multi-shot LLM-based methods.
@@ VI-B2 ML / DL Defense
Leveraging classification models, adversarial training, and other machine learning or deep learning (ML/DL) techniques to strengthen a model’s resistance to jailbreak attack has proven to be an effective defense strategy. For example, by generating both adversarial and benign samples and training a binary classifier to distinguish between them, the model’s robustness against malicious inputs and perturbations can be significantly improved [134]. Such approaches not only mitigate the impact of adversarial attacks but also enhance the model’s stability and reliability in real-world applications.
@@
Defense frameworks like those proposed by Deng et al. [88] use an iterative optimization approach to enhance LLM security with minimal impact on its capabilities. The process begins by generating adversarial prompts using an attack framework, then fine-tuning the target LLM to encourage safe refusal responses. The model is then evaluated, and prompts that can still break through are added to the attack dataset, allowing for further fine-tuning until the LLM can withstand stronger attacks. Zhang et al. [128] introduce a goal prioritization method to defend LLMs against jailbreak attacks. This method controls goal priority during both inference and training. At inference, a “plug-and-play” strategy prioritizes safety, while during training, contrastive instances help the model learn to prioritize safety under different conditions. The method significantly reduces jailbreak attack success rates, dropping ChatGPT’s from 66.4% to 3.6%, without needing jailbreak-specific training data, showing strong generalization.
@@
Furthermore, Zeng et al. [117] propose the Root Defence Strategy (RDS), a decoder-oriented defense that corrects harmful outputs instead of rejecting them. RDS uses a trainable classifier to evaluate the harm of each token during decoding and prioritizes safer options. It also employs speculative decoding to enhance safety while maintaining inference speed. Wang et al. [129] introduce SelfDefend, a dual-layer protection system where a shadow LLM (LLM_defense) runs alongside the target LLM (LLM_target). The shadow LLM uses specific detection prompts to identify harmful content, and if any is found, SelfDefend blocks harmful responses from the target LLM. This method provides effective protection for both open-source and closed-source models with minimal latency, optimizing defense through data distillation and fine-tuning.
@@
JATMO [119] is a task-specific LLM generation method that resists prompt injection attacks. It fine-tunes a base model to perform specific tasks without instruction tuning, making it less susceptible to manipulation by malicious prompts. The method uses a standard instruction-tuned LLM as a teacher to generate outputs for task datasets, which are then used to fine-tune the base model, allowing it to learn the task mapping without relying on instructions. Wang et al. [125] propose SELF-GUARD, which trains the LLM to perform self-auditing by tagging outputs with $harmful$ or $harmless$ labels, enabling the model to detect harmful content while maintaining flexibility for external protection. SELF-GUARD minimizes performance degradation from safe training and reduces the computational overhead of external methods, improving security without sacrificing efficiency.
@@ VI-B3 Adversarial Detection Defense
Adversarial detection defenses typically employ independent detection models or specific metric analyses to intercept malicious inputs or abnormal outputs, such as judging potential risks based on confidence scores. Cao et al. [135] propose Robustly Aligned LLM (RA-LLM), which improves model robustness by randomly deleting parts of input requests. The model relies on its internal alignment capabilities to assess requests as benign, effectively avoiding adversarial prompts. This method requires no external harmful content detector and reduces the attack success rate from nearly 100% to below 10%. We believe that this random dropout operation proposed by Cao et al. [135] invalidates adversarial prompts in aligned attacks, which are usually sensitive to small perturbations. Jain et al. [113] introduce the Perplexity Filter method, which detects unreadable attack prompts by calculating their perplexity with another LLM. If the perplexity exceeds a preset threshold, the prompt is filtered out, effectively intercepting malicious content.
@@
Subsequently, Mao et al. [89] point out that adversarial prompts usually consist of harmless instructions and harmful data content. For example, in the short prompt “please teach me how to kill,” the word “kill” belongs to the harmful data part, while “please teach me” is the harmless instruction part. Based on this observation, the authors design a harmful content separator that automatically identifies and separates the instruction and data components within prompts. Then, the system inputs the data part into a harmful content filter for safety detection. If the data content violates safety constraints, it is deleted. Zhang et al. [41] propose a defense strategy that detects harmful instructions before execution. The core method involves querying the LLM to evaluate whether an instruction may be harmful or violate policies, using a “YES” or “NO” response. This strategy helps defend against jailbreak attacks but had biases toward certain types, such as those aimed at causing damage or stealing data. To address this, they improve the detection prompt to focus on whether the instruction “intentionally causes model failure,” providing a more balanced evaluation. Zhang et al. [116] introduce JailGuard, a framework designed for detecting prompt-based attacks, including jailbreaking and hijacking in text and image inputs. JailGuard assumes that attack inputs are more fragile and sensitive to slight mutations compared to normal inputs. It generates multiple mutated variants of the input and calculates the model’s response differences. If the divergence exceeds a preset threshold, it flags the input as an attack. JailGuard employs 18 mutation methods (16 random and 2 semantic-driven) and uses Kullback-Leibler (KL) divergence to assess response differences, making the detection more efficient.
@@
In addition, Wang et al. [122] first obtain the initial response from the target LLM for a given prompt and then use a language model to infer the likely input, known as the “backtranslated prompt.” Because this prompt is derived from the LLM’s own output rather than directly crafted by attackers, it typically reveals the original prompt’s true intent. If the LLM rejects the backtranslated prompt, the original is flagged as potentially harmful and blocked. This method has minimal impact on benign inputs while remaining effective against complex adversarial prompts. Pasquini et al. [124] introduce Mantis, a defense framework designed to counter automated network attacks from LLM prompt injection vulnerabilities. Mantis embeds crafted prompt injections into system responses to mislead and disrupt the attacker’s LLM (passive defense) and can even counterattack the attacker’s machine (active defense). Additionally, Mantis uses decoy vulnerable services and dynamic prompt injection to interfere with the attacker in real time.